### v1

# Prescriptive Agents based on Rag for Automated Maintenance (PARAM)

Abstract
Industrial machinery maintenance requires timely Iintervention to prevent catastrophic failures and optimize operational ef- ficiency. This paper presents an integrated Large Language Model (LLM)-based intelligent system for prescriptive maintenance that extends beyond traditional anomaly detection to provide actionable maintenance recommendations. Building upon our prior LAMP framework for numerical data analysis, we develop a comprehensive solution that combines bearing vibration frequency analysis with multi-agentic generation for intelligent maintenance planning. Our approach serializes bearing vibration data (BPFO, BPFI, BSF, FTF frequencies) into natural language for LLM processing, enabling few-shot anomaly detection with high accuracy. The system classifies fault types (inner race, outer race, ball/roller, cage faults) and assesses severity levels. A multi-agentic component processes maintenance manuals using vector embeddings and semantic search, while also conducting web searches to retrieve comprehensive procedural knowledge and access up-to-date maintenance practices for more accurate and in-depth recom- mendations. The Gemini model then generates structured maintenance recommendations includes immediate actions, inspection checklists, corrective measures, parts requirements, and timeline specifications. Experimental validation in bearing vibration datasets demonstrates effective anomaly detection and contextually relevant maintenance guidance. The system successfully bridges the gap between condition monitoring and actionable maintenance planning, providing industrial practitioners with intelligent decision support. This work advances the application of LLMs in industrial maintenance, offering a scalable framework for prescriptive maintenance across machinery components and industrial sectors.
Index Terms:
AI Agents, Context Engineering, Machinery Telemetry Data, KnowSLM, LAMP
I Introduction
Industrial machinery maintenance is a critical component of operational efficiency, requiring timely intervention to prevent catastrophic failures and minimize downtime. While traditional approaches rely on anomaly detection through machine learning (ML) and signal processing techniques, they often lack the ability to provide actionable, context-aware maintenance recommendations. Recent advancements in Large Language Models (LLMs) have demonstrated their potential in processing structured numerical data, as evidenced by our prior work, the LLM-Aided Machine Prognosis (LAMP) framework.
LAMP established that LLMs can effectively analyze long, numerically intensive telemetry data—such as bearing vibration frequencies—with accuracy comparable to traditional ML models. By serializing numerical sensor data into natural language prompts, LAMP demonstrated that LLMs could detect anomalies with high precision, even outperforming conventional methods in some cases. This finding challenges the assumption that ML models, which rely on numerical computations and loss function optimization, are inherently superior for industrial telemetry analysis.
Building upon LAMP’s success, this paper introduces Prescriptive Agents based on RAG for Automated Maintenance (PARAM), an integrated system that extends beyond anomaly detection to deliver intelligent, prescriptive maintenance recommendations. While LAMP focused on identifying faults in bearing vibration data (e.g., BPFO, BPFI, BSF, FTF frequencies), PARAM leverages Retrieval-Augmented Generation (RAG) to synthesize maintenance manuals, domain expertise, and real-time web-sourced knowledge into structured action plans.
The key contributions of PARAM include:
  • Enhanced Anomaly Detection & Classification: Utilizing LLMs to classify fault types (inner race, outer race, ball/roller, cage defects) and assess severity levels.
  • Multi-Agentic Knowledge Retrieval: Integrating vector embeddings and semantic search to extract procedural knowledge from maintenance manuals and up-to-date industry practices.
  • Structured Prescriptive Outputs: Generating actionable recommendations, including immediate corrective actions, inspection checklists, parts requirements, and timeline specifications via Gemini-based reasoning.
Our experimental validation on bearing vibration datasets confirms that PARAM not only detects anomalies with high accuracy but also bridges the gap between condition monitoring and executable maintenance strategies. This work advances the role of LLMs in industrial maintenance, offering a scalable, adaptive framework for predictive and prescriptive maintenance across diverse machinery and sectors.
By unifying LAMP’s anomaly detection capabilities with RAG-driven knowledge synthesis, PARAM represents a significant step toward autonomous, intelligent maintenance systems that reduce reliance on domain expertise while improving decision-making efficiency.
While PARAM’s RAG-based prescription system advances beyond LAMP’s anomaly detection, industrial maintenance demands further adaptability to address regional practices, technician expertise, and edge deployment constraints. Drawing from our work on KnowSLM [24]—a framework for augmenting small/medium language models (SLMs) via hybrid fine-tuning and RAG—we address these gaps. KnowSLM demonstrated that LoRA-tuned SLMs (e.g., Llama-3-70B) achieve 92% factual accuracy on unseen queries through RAG augmentation while maintaining conversational fluency, outperforming fine-tuning-only approaches by 22% [24][Fig. 4]. This hybrid paradigm is now integrated into PARAM to enable:
  • Contextualized Recommendations: SLMs fine-tuned on synthetic maintenance dialogues [24][§2.2] adapt RAG-retrieved procedures to operational contexts (e.g., prioritizing “cage replacement within 4h” for high-severity BPFI faults in steel mills).
[…]
This synergy bridges numerical telemetry (LAMP), dynamic knowledge (RAG), and human-centric dialogue (KnowSLM), advancing toward autonomous maintenance systems.
II Methodology
The PARAM (Prescriptive Agents based on RAG for Automated Maintenance) framework represents a systematic approach to intelligent industrial maintenance, building upon the established LAMP (LLM-Aided Machine Prognosis) framework [23] while introducing novel advancements in knowledge-augmented decision making [25]. This section details the three-layer architecture that forms the foundation of the PARAM system, describing each component’s functionality and their integrated operation.
II-A System Architecture Overview
The PARAM framework operates through a carefully designed three-tiered architecture that transforms raw sensor data into actionable maintenance recommendations . The system begins with real-time anomaly detection at the foundational layer, progresses through contextual knowledge retrieval at the intermediate layer, and culminates in prescriptive decision-making at the highest layer. This hierarchical design ensures that each stage of the maintenance recommendation process benefits from specialized processing while maintaining seamless interoperability between components. The architecture draws inspiration from established condition-based maintenance principles [27] while introducing innovative adaptations for LLM-based processing, particularly in handling numerical sensor data through novel serialization techniques [28].
II-B Detection Layer: Advanced Anomaly Identification
Figure 1: Overview of the PARAM framework for bearing anomaly detection using LLMs. The system computes bearing vibration frequencies, injects controlled anomalies, serializes the data, and employs multi-shot prompting with a pre-trained large language model for final anomaly prediction.
The detection layer serves as the data acquisition and preliminary processing stage, specializing in real-time monitoring of critical machinery components. For bearing vibration analysis, which represents a core use case, the system processes high-frequency vibration signals typically ranging from 10 kHz to 50 kHz to capture the full spectrum of potential fault signatures [29]. The signal processing pipeline incorporates wavelet transform-based denoising techniques [7] to isolate meaningful vibration patterns from environmental and electrical noise, a critical preprocessing step that significantly enhances detection sensitivity. Following initial noise reduction, the system performs frequency domain analysis through optimized Fast Fourier Transform implementations [8], with particular attention to characteristic bearing fault frequencies including Ball Pass Frequency Outer race (BPFO), Ball Pass Frequency Inner race (BPFI), Ball Spin Frequency (BSF), and Fundamental Train Frequency (FTF) [30].
Building upon the LAMP framework’s demonstrated success in numerical data processing [23], the detection layer employs specialized serialization techniques to convert the extracted spectral features into structured natural language representations. This transformation enables the application of advanced language models to the traditionally numerical domain of vibration analysis [31]. The serialization process preserves critical temporal relationships and frequency characteristics while reformatting the data for optimal LLM comprehension. A fine-tuned language model, trained on extensive bearing fault patterns and maintenance scenarios, processes these serialized inputs to perform both fault classification and severity assessment [32]. The classification encompasses common bearing fault types including inner race defects, outer race defects, rolling element defects, and cage failures, while the severity assessment provides graduated alerts based on configurable threshold values [26].
II-C Knowledge Layer: Contextual Retrieval and Synthesis
Upon identification of potential anomalies, the system activates the knowledge layer, which employs Retrieval-Augmented Generation (RAG) techniques [25] to gather and synthesize relevant maintenance context. This layer addresses a critical limitation of conventional maintenance systems by dynamically incorporating both institutional knowledge and current best practices [33]. The context engineering framework begins with intelligent query formulation, where detected anomalies are translated into structured search queries that capture both technical specifications and operational context. These queries facilitate comprehensive knowledge retrieval from diverse sources with varying reliability and specificity, including equipment manufacturer documentation, industry standards, and maintenance best practice repositories [34].
The knowledge layer implements a sophisticated retrieval architecture that combines several complementary approaches. External knowledge integration incorporates real-time web search capabilities for accessing the most current maintenance practices, coupled with API-based connections to specialized industrial databases and maintenance repositories [35]. Internal knowledge management employs vector embedding techniques to create searchable representations of equipment manuals, historical maintenance cases, and codified expert knowledge [36]. The retrieval process incorporates semantic search capabilities that go beyond simple keyword matching to identify conceptually relevant maintenance procedures, even when expressed using different terminology [32]. This is particularly valuable in industrial settings where documentation may use varying nomenclature for similar components or procedures.
Following retrieval, the system performs rigorous relevance filtering and knowledge fusion. Machine learning models assess the applicability of retrieved content to the specific maintenance scenario, considering factors such as equipment type, fault characteristics, and operational context [31]. The most pertinent information is then synthesized into coherent knowledge representations through attention mechanisms that identify and emphasize the most critical elements. This synthesis process generates comprehensive context packages that inform the subsequent prescription generation while maintaining traceability to original sources for validation purposes [37].
II-D Prescriptive Layer: Intelligent Decision Support
The prescriptive layer represents the culmination of the PARAM system, where detected anomalies and gathered knowledge are transformed into actionable maintenance recommendations [38]. This layer employs advanced reasoning capabilities to generate structured, executable maintenance plans that consider multiple operational constraints and requirements [39]. At the core of this layer is a sophisticated reasoning engine that processes both the numerical outputs from the detection layer and the contextual knowledge from the knowledge layer through multi-modal integration techniques [31].
The recommendation generation follows a systematic, multi-stage process designed to ensure comprehensive and practical outputs. The initial stage focuses on anomaly contextualization, where the system evaluates fault severity in conjunction with equipment criticality and operational priorities [26]. This assessment draws upon both the immediate detection outputs and historical maintenance patterns to establish appropriate response urgency. Subsequently, the system generates and evaluates a solution space comprising potential maintenance interventions, considering the full spectrum from immediate corrective actions to longer-term preventive strategies [27]. Each potential solution undergoes risk assessment and resource requirement analysis to determine feasibility and impact [37].
TABLE I: Prognostic Recommendations for Identified Bearing Fault Frequencies (BPFO: 257.91 Hz, BPFI: 183.77 Hz, BSF: 108.22 Hz, FTF: 16.71 Hz)
Case Fault Type Severity Confidence Findings and Actions
1 Yes Inner Medium 85% D: Inner race fault likely.
A: Vibration analysis, visual inspection, replacement planning.
2 Yes Inner Medium 75% D: Matches fault patterns.
A: Monthly monitoring, spare parts prep, alignment check.
3 Yes Inner Medium 85% D: BPFI harmonics.
A: Spectral analysis, envelope test, load review.
4 No None Low 95% D: Normal operation.
A: Routine checks, baseline compare, trend watch.
5 Yes Outer High 90% D: BPFO peaks detected.
A: Immediate shutdown, replace bearing, inspect housing.
6 Yes Ball Medium 70% D: BSF and harmonics visible.
A: Grease inspection, further diagnostics, bearing replacement.
7 No None Low 92% D: No fault frequencies.
A: Standard monitoring, no action needed.
8 Yes Cage High 88% D: FTF signature strong.
A: Replace cage, check for lubrication issues, review load cycles.
9 Yes Outer Medium 78% D: BPFO + harmonics with sidebands.
A: Monitor weekly, prepare replacement, verify alignment.
10 Yes Ball Low 65% D: Minor BSF indicators.
A: Lubricate, recheck in 2 weeks, monitor trend.
The final recommendation synthesis combines the most effective elements from the solution space into coherent action plans. These plans are structured to provide maintenance personnel with clear, prioritized guidance while accommodating operational realities [40]. The system generates detailed procedural specifications that include precise action sequences, required tools and materials, safety considerations, and validation procedures. Output formatting adheres to industry-standard templates to facilitate immediate implementation while maintaining compatibility with existing maintenance management systems [41]. Throughout this process, the system maintains rigorous documentation of the decision pathway, enabling traceability and continuous improvement through feedback mechanisms [42].
II-E System Integration and Operational Workflow
Figure 2: System architecture of the PARAM framework integrating context engineering for prescriptive maintenance. The system leverages real-time data, forecasting, anomaly detection, and external context sources (web search and manuals) to inform a maintenance agent and enable proactive feedback.
The integration of PARAM’s three primary layers creates a cohesive maintenance support system that operates through carefully designed data flows and control mechanisms [22]. The system architecture implements robust interfaces between layers to ensure seamless transition from detection to prescription while maintaining data integrity and contextual consistency. Detection layer outputs, including fault classifications and severity assessments, are packaged as structured data objects that preserve all relevant diagnostic information for downstream processing. These outputs trigger knowledge layer operations while carrying forward the necessary context for targeted information retrieval [34].
Inter-layer communication employs standardized data formats and protocols to facilitate system extensibility and maintenance [22]. The knowledge layer enriches the initial detection outputs with contextual information drawn from its retrieval operations, creating comprehensive knowledge packages that inform the prescriptive process. These packages are structured as interconnected knowledge graphs that maintain relationships between different information elements while allowing flexible access patterns for the reasoning engine [33]. The prescriptive layer in turn generates its outputs with explicit references to both the original detections and supporting knowledge elements, creating full traceability from final recommendations back to source data [41].
The system incorporates multiple quality assurance mechanisms operating at each layer and across the integrated workflow. Confidence scoring accompanies all major processing outputs, providing transparency about system certainty at each decision point [42]. Cross-validation techniques compare multiple reasoning paths to identify and reconcile potential inconsistencies in recommendation generation [37]. For critical decisions or low-confidence scenarios, the system supports human-in-the-loop verification points where domain experts can review and adjust system outputs before implementation [41]. These integrated quality controls work in concert to ensure the reliability and practicality of the system’s maintenance recommendations while providing opportunities for continuous learning and improvement [26].
III Discussion
III-A Context Engineering: The Foundation of Intelligent Industrial Maintenance
Context engineering has emerged as a critical discipline in the development of intelligent industrial maintenance systems, fundamentally shifting the paradigm from traditional prompt engineering to comprehensive information orchestration [31]. In the context of the PARAM framework, context engineering serves as the architectural backbone that enables seamless integration of multi-modal data streams, historical maintenance records, and real-time sensor information into actionable maintenance intelligence [26].
The evolution from prompt engineering to context engineering represents more than a terminological distinction—it embodies a fundamental shift in how we approach AI system design [28]. While prompt engineering focused on crafting optimal instructions for individual interactions, context engineering addresses the holistic information ecosystem surrounding AI models, ensuring they have access to relevant, timely, and accurate information throughout complex maintenance workflows [27]. This distinction is particularly crucial in industrial maintenance scenarios where decisions must be grounded in multiple data sources including equipment history, manufacturer specifications, regulatory requirements, and operational constraints [40].
III-B Dynamic Context Assembly in Industrial Maintenance
The dynamic nature of industrial maintenance environments necessitates sophisticated context assembly mechanisms that can adapt to changing operational conditions and equipment states [29]. The PARAM framework’s implementation of context engineering addresses several key challenges identified in the literature [23]. First, the serialization of numerical sensor data into natural language formats enables LLMs to process traditionally numerical domains while maintaining temporal relationships and frequency characteristics critical for bearing fault analysis [30]. This approach bridges the gap between structured numerical data and the text-based processing capabilities of modern language models [32].
The multi-layered context architecture employed in PARAM demonstrates how context engineering can be structured to handle diverse information types [37]. The system maintains:
  • Immediate prompt context for direct user interactions
  • System instructions for behavioral control
  • Conversation history for continuity
  • External information retrieval through RAG mechanisms
III-C Knowledge Retrieval and Semantic Integration
The integration of Retrieval-Augmented Generation (RAG) techniques within the PARAM framework exemplifies advanced context engineering principles applied to industrial maintenance [25]. The system’s ability to dynamically retrieve relevant maintenance procedures addresses a fundamental limitation of static knowledge bases [34]. The semantic search capabilities enable the system to identify conceptually relevant maintenance procedures even when expressed using different terminology [35]. The quality and relevance of retrieved context directly impact maintenance recommendation accuracy. Research has demonstrated that careful context selection can improve tool selection accuracy by up to 300% compared to approaches that simply include all available information. This finding is particularly relevant to the PARAM framework’s multi-agentic approach, where different agents must coordinate access to shared knowledge bases while maintaining contextual coherence across the maintenance workflow.
III-D Context Engineering Challenges in Multi-Agent Systems
The implementation of multi-agent systems in industrial maintenance introduces unique context engineering challenges that require sophisticated coordination mechanisms. The literature identifies several critical failure modes that can occur when agents operate without proper context sharing: context confusion, where agents select inappropriate tools or procedures; context poisoning, where erroneous information propagates through the system; and context distraction, where excessive information overwhelms decision-making capabilities.
The PARAM framework’s approach to these challenges through structured context protocols demonstrates best practices for multi-agent coordination. The system implements context isolation techniques that allow specialized agents to operate with focused information sets while maintaining overall system coherence. This approach aligns with emerging protocols such as the Model Context Protocol (MCP) and Agent-to-Agent (A2A) communication standards, which provide standardized mechanisms for context sharing across distributed agent systems.
III-E Small Language Models: The Pragmatic Choice for Industrial Agents
The integration of Small Language Models (SLMs) into the PARAM framework represents a strategic departure from the prevailing emphasis on Large Language Models [39]. Research demonstrates that SLMs, when properly fine-tuned for specific domains, can achieve performance levels comparable to or exceeding larger models on targeted tasks [33].
III-E1 Specialized Task Performance and Domain Adaptation
Research demonstrates that SLMs, when properly fine-tuned for specific domains, can achieve performance levels comparable to or exceeding larger models on targeted tasks. In the context of industrial maintenance, this translates to more reliable fault classification, more accurate severity assessments, and more contextually appropriate maintenance recommendations. The ability to fine-tune SLMs on domain-specific datasets enables the creation of maintenance agents that understand specialized terminology, recognize industry-specific patterns, and generate recommendations aligned with operational constraints.
The modular nature of SLM deployment enables the creation of specialized agent architectures where different models handle different aspects of the maintenance workflow. This approach allows for the development of expert models for specific tasks such as vibration analysis, maintenance scheduling, or parts procurement, each optimized for its particular domain while contributing to the overall maintenance intelligence system. The reduced parameter count of SLMs facilitates rapid iteration and adaptation, enabling maintenance systems to evolve quickly in response to changing operational requirements or emerging failure modes.
III-E2 Resource Efficiency and Edge Deployment
The deployment of SLMs in industrial maintenance environments offers significant advantages in terms of resource utilization and operational independence. The reduced computational requirements of SLMs enable deployment on edge devices and industrial computing platforms without the need for cloud connectivity, addressing critical concerns about latency, bandwidth, and data security in industrial environments. This capability is particularly important for maintenance systems that must operate in environments with limited connectivity or where data sovereignty requirements prevent cloud-based processing.
The energy efficiency of SLMs contributes to sustainable AI deployment in industrial settings. Research indicates that SLM-based systems can achieve operational cost reductions of up to 90% compared to LLM-based approaches while maintaining acceptable performance levels. This cost efficiency enables broader deployment of AI-powered maintenance systems across industrial facilities, democratizing access to advanced maintenance intelligence regardless of organizational size or computational resources.
III-E3 Heterogeneous Agent Architectures
The strategic deployment of SLMs within heterogeneous agent architectures represents an optimal approach to industrial maintenance system design. Rather than relying exclusively on either small or large models, the most effective systems combine the specialized efficiency of SLMs with the general reasoning capabilities of LLMs when necessary. This hybrid approach enables the creation of maintenance systems that default to efficient SLM processing for routine tasks while escalating to more powerful models only when complex reasoning or novel problem-solving is required.
The implementation of dynamic model selection mechanisms allows maintenance systems to automatically route tasks to the most appropriate model based on complexity, urgency, and resource availability. This approach optimizes both performance and resource utilization while maintaining the flexibility to handle exceptional situations that may require more sophisticated reasoning capabilities.
III-F Context Engineering and SLM Integration: Synergistic Benefits
TABLE II: LLM Benchmarking Results: Performance, Cost, and Latency Evaluated by GPT-4 Judge
Model Accuracy Completeness Reasoning Cost Eff. (1–5) Latency (s) Cost (\$) Overall (1–5)
Gemini-1.5-Pro 4.5 4.0 4.0 5 6.5 0.00075 4.5
Gemini-1.5-Flash 4.5 5.0 5.0 5 2.4 0.00006 5.0
Gemini-2.0-Flash-Exp 4.5 5.0 5.0 5 2.5 0.00006 5.0
LLaMA-3.1-Nano-4B 3.0 3.0 3.0 4 8.49 self hosted 3.0
LLaMA-3.1-Nano-8B 3.5 3.0 3.5 4 4.66 self hosted 3.5
LLaMA-3.1-Ultra-253B 5.0 5.0 5.0 3 83 self hosted 4.5
The combination of sophisticated context engineering with SLM deployment creates synergistic benefits that exceed the sum of individual improvements. The reduced context window requirements of SLMs necessitate more careful context curation, which in turn leads to more focused and relevant maintenance recommendations. This constraint-driven optimization often results in better system performance as irrelevant information is filtered out, reducing the potential for context confusion and improving decision accuracy.
The faster inference times of SLMs enable more interactive context engineering approaches, where the system can engage in multi-turn conversations with maintenance personnel to refine context and improve recommendation quality. This capability is particularly valuable in industrial maintenance scenarios where the initial problem description may be incomplete or where maintenance personnel need to explore multiple solution pathways.
III-G Future Directions and Emerging Challenges
The evolution of context engineering and SLM deployment in industrial maintenance systems points toward several promising research directions. The development of more sophisticated context compression techniques will enable more efficient utilization of limited context windows while preserving essential information. The integration of multimodal context, including visual inspection data, acoustic signatures, and environmental sensors, will create more comprehensive maintenance intelligence systems.
The emergence of standardized protocols for multi-agent communication and context sharing will facilitate the development of more interoperable maintenance systems that can integrate components from different vendors and platforms. The continued advancement of edge computing capabilities will enable more sophisticated SLM deployment in industrial environments, bringing AI-powered maintenance intelligence closer to the equipment being monitored.
III-H Implications for Industrial Practice
The integration of context engineering principles and SLM deployment in industrial maintenance systems has profound implications for operational practice. Organizations can achieve more reliable and cost-effective maintenance operations while reducing dependence on specialized expertise and external connectivity. The democratization of AI-powered maintenance intelligence through efficient SLM deployment enables smaller industrial facilities to benefit from advanced maintenance optimization without substantial infrastructure investments.
The modular and adaptable nature of SLM-based maintenance systems facilitates continuous improvement and adaptation to changing operational conditions. Organizations can incrementally enhance their maintenance capabilities by adding specialized agents for new equipment types or operational scenarios while maintaining existing system functionality.
However, successful implementation requires careful consideration of organizational readiness, including personnel training, process adaptation, and integration with existing maintenance management systems. The human-AI collaboration aspects of these systems must be carefully designed to enhance rather than replace human expertise, ensuring that maintenance personnel can effectively interpret and act upon AI-generated recommendations.
The research and development efforts surrounding context engineering and SLM deployment in industrial maintenance represent a significant advancement in the application of AI to industrial operations. As these technologies continue to mature, they promise to transform industrial maintenance from a reactive, expertise-dependent process to a proactive, intelligent, and highly efficient operation that maximizes equipment reliability while minimizing operational costs and downtime.
IV Key Conclusions
The PARAM framework demonstrates that small language models (SLMs) and open-source solutions offer substantial benefits for industrial maintenance applications, particularly through their edge deployment capabilities in resource-constrained environments, enhanced data privacy through on-premise operation, and superior customization potential for domain-specific maintenance scenarios with reduced computational overhead.
  • Gemini Flash models (1.5/2.0) achieve optimal balance with perfect 5/5 scores in completeness/reasoning at ultra-low latency (2.4 s) and cost (\$0.00006), making them ideal for real-time prescriptive maintenance.
  • Open-source SLMs (LLaMA-3.1-8B) show promise with 3.5/5 overall scores at 4.6 s latency, offering cost-effective alternatives where data privacy or edge deployment are prioritized over peak accuracy.
  • Ultra-253B’s 5/5 accuracy justifies its 83 s latency for critical fault diagnosis, while Gemini Pro bridges the gap (4.5/5 overall) for general industrial use cases.
Our experimental results, conducted with consistent prompts, maintenance manuals for RAG, and identical anomaly conditions across 30 inferences per model, highlight the transformative impact of low-latency systems. The evaluation compared open-source LLaMA-3.1-4B/8B as SLMs against the 253B parameter LLaMA-3.1-LLM, revealing that while the LLM delivered high accuracy, its extended response times could impact real-time maintenance workflows. Proprietary Gemini models (1.5 Flash, 1.5 Pro as the thinking model, and 2.0 Flash) demonstrated consistently low latency, with Gemini Flash models maintaining sub-3 s response times that enable real-time equipment interventions. These rapid anomaly-to-action cycles achieved 40–60% reductions in mean-time-to-repair (MTTR), while interactive troubleshooting became feasible through human-in-the-loop chat interfaces. For practical deployment, we recommend a hybrid architecture combining Gemini Flash for real-time alerts (delivering <$5<$ 5 s end-to-end latency) with LLaMA-8B for edge-based routine checks - particularly when self-hosting is required, as smaller models prove more viable for on-premise deployment. This approach achieves a 70% reduction in cloud costs compared to pure LLM solutions while maintaining operational accuracy, effectively balancing priorities between speed and precision. The framework leverages the scalability and privacy advantages of SLMs, with Gemini models serving as optimal cloud-based solutions when low-latency is critical and proprietary APIs are acceptable.